Claire, who is six years older then her sister Amy, isnow three times as old as Amy was four years ago. How old are Claire and Amy now?
Answer by nerdybill(7008)

.
Let x = Amy's age
then from "Claire, who is six years older then her sister Amy"
x+6 = Claire's age
.
From:" is now three times as old as Amy was four years ago."
x+6 = 3(x-4)
x+6 = 3x-12
6 = 2x-12
18 = 2x
9 years = x (Amy's age)
.
Claire's age:
x+6 = 9+6 = 15 years